OpenStudy (bbb911):

2+ 1/6 ------- 1- 4/3

OpenStudy (nubeer):

can you use calculator?

OpenStudy (bbb911):

nope

OpenStudy (anonymous):

={(12/6)+(1/6)}/{(3/3)-(4/3)} =(13/6)/(-1/3) =(13/6)*(-3) =-13/2

OpenStudy (bbb911):

thts wat i got thnks :)

OpenStudy (anonymous):

anytime :)